Suppose the production function is given by Q(L,K) = L 4/3

K 4

1

. Assuming capital is fixed,

find APL and MPL.


1
Expert's answer
2021-08-06T08:15:50-0400

"AP_L=\frac {TP}{L}"

"AP_L=\frac {L^\frac{4}{3}K^\frac{1}{4}}{L}=L^\frac{1}{3}K^\frac{1}{4}"

"MP_L=\frac{\Delta TP_L}{\Delta L}"

"MP_L=\frac{4}{3}L^\frac{1}{3}K^\frac{1}{4}"